# Is the Faber-Krahn inequality true for the Stokes operator?

### v1

@@ 1.1 Setting
The optimisation of spectral quantities with respect to a domain is a central question in shape optimisation and in the calculus of variations. Of particular importance is the paradigmatic question of minimising (with respect to the domain ${\Omega}$) the first eigenvalue of a differential operator under a volume constraint on ${\Omega}$. The celebrated Faber-Krahn inequality asserts that, when the operator is the (scalar) Dirichlet-Laplacian, the ball minimises the first eigenvalue with a volume constraint In the present paper, we investigate the minimisation of the first eigenvalue of the Dirichlet-Stokes operator. Namely, for a bounded open set $\Omega$ consider the first eigenvalue
$\lambda_{1}({\Omega}):=\min_{\begin{subarray}{c}{\mathbf{u}}\in W^{1,2}_{0}({\Omega})^{3}\\ \nabla\cdot{\mathbf{u}}=0\text{ in }{\Omega},\ {\mathbf{u}}\neq 0\end{subarray}}\frac{\int_{\Omega}\|{\nabla}{\mathbf{u}}\|^{2}}{\int_{\Omega}\|{\mathbf{u}}\|^{2}},$ (1)
@@
$\begin{cases}-\Delta{\mathbf{u}}+{\nabla}p=\lambda_{1}({\Omega}){\mathbf{u}}&\text{ in }{\Omega}\,,\\ {\nabla}\cdot{\mathbf{u}}=0&\text{ in }{\Omega}\,,\\ {\mathbf{u}}=0&\text{ on }{\Omega}.\end{cases}$ (2)
In (2) the function $p$ (unique up to an additive constant) is the pressure, which can be interpreted as the Lagrange multiplier associated with the incompressibility constraint.
@@
The constant $V_{0}>0$ is a given volume constraint (which will be immaterial as we will be led to work with a scale invariant formulation of (3)).
@@
    First, we obtain an existence result in the class of quasi-open sets (this is the natural framework for such existence results; we refer to Definition 1) This is Theorem 1. The method of proof relies on the concentration-compactness principle of Lions [37], which was adapted by Bucur [3] to the setting of shape optimisation. The main difficulty here is the incompressibility condition.
@@
We discuss related works in details in section 1.3 but let us already highlight some aspects of our problem. One of the first results in spectral shape optimisation is the celebrated Faber-Krahn inequality [23, 33], which asserts that, under a volume constraint, the ball minimises the first eigenvalue of the Dirichlet-Laplacian in any dimension. This inequality can be derived from numerous methods, some of which can be generalised to the minimisation of other spectral quantities and the ball is very often the minimiser (or at least a local minimiser) for the first eigenvalue of several scalar differential operators. Thus, Theorem 3 hints at a deeper discrepancy between the scalar and the vectorial case. In general, let us observe that the literature devoted to spectral optimisation problems in the vectorial case is scarce. To the best of our knowledge, the results closest to ours were derived very recently by Enciso, Gerner & Peraltas-Sala [22, 21, 25] in the case of the ${\mathbf{curl}}$ operator, a problem which was also investigated by Cantarella, DeTurck, Gluck & Teytel in the early 2000’s [11].
@@
    For any $k\in\mathbb{N}\,,p\in[1;+\infty)$, $W^{k,p}({\Omega})$ (resp. $W^{k,p}_{0}({\Omega})$) denotes the usual Sobolev space of order $k$ and index $p$ (resp. the functions of $W^{k,p}({\Omega})$ whose trace on $\partial{\Omega}$ is zero).
@@
    The double-dot product of two matrices $A=(a_{ij})_{1\leqslant i,j\leqslant N}$ and $B=(b_{ij})_{1\leqslant i,j\leqslant N}$ is the real number $A:B$ given by $A:B=\sum_{i,j=1}^{N}a_{ij}b_{ij}$.
@@
    $\mathbb{B}_{d}=\{x\in\mathbb{R}^{d}\ |\ \|x\|<1\}$, the Euclidean unit ball of $\mathbb{R}^{d}$.
@@
  8. 8.
    $\operatorname{Per}(\Omega)$, for $\Omega$ bounded connected, denotes the perimeter of $\Omega$ in the sense of De Giorgi
@@
To set our problem in an appropriate (from the point of view of existence properties) framework, we recall the definition of quasi-open sets [30, Chapter 3]:
@@
$\mathrm{cap}(\omega):=\sup_{\begin{subarray}{c}K\textnormal{compact}\\ K\subset\omega\end{subarray}}\,\inf_{\begin{subarray}{c}v\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{d})\\ v\geqslant 1\text{ on }K\end{subarray}}\int_{K}\left|{\nabla}v\right|^{2}.$
@@
Quasi-open sets are the natural framework in which to consider spectral shape optimisation problems [10, 49]. For any given ${\Omega}\in\mathscr{O}$ we can define [30, Chapter 3] the Sobolev space $W^{1,2}_{0}({\Omega})$ as the set of functions $v\in W^{1,2}(\mathbb{R}^{d})$ that are equal to 0 on $\mathbb{R}^{d}\backslash{\Omega}$ up to a set of zero capacity. For any ${\Omega}\in\mathscr{O}$ we set
$\lambda_{1}({\Omega}):=\min_{\begin{subarray}{c}{\mathbf{u}}\in W^{1,2}_{0}({\Omega})^{3}\\ \nabla\cdot{\mathbf{u}}=0\text{ in }{\Omega}\,,{\mathbf{u}}\neq 0\end{subarray}}\frac{\int_{\Omega}\|{\nabla}{\mathbf{u}}\|^{2}}{\int_{\Omega}\|{\mathbf{u}}\|^{2}}.$ (4)
@@ 1.2 Main results of the paper
The proof of this theorem relies, in this vectorial setting, of the approach to the Faber-Krahn inequality of Bucur & Freitas [5, Proposition 3.1] or Bucur & Varchon [6], which in turn relies on the concentration-compactness principle of Lions [37].
Of course, if $\mathscr{O}$ were to be replaced with $\mathscr{O}_{D}:=\{{\Omega}\in\mathscr{O}\,,\mathscr{O}\subset D\}$ for a given compact set $D$ (“box constraint”), the existence of a minimiser would follow from an adaptation of the Buttazzo-Dal Maso theorem [10], although one should be cautious when handling the zero-divergence constraint. This fact is actually used in the proof of our existence result. We refer to section 1.3 for more references on shape optimisation without box constraints.
@@
When we work in the two-dimensional case, if the domain ${\Omega}$ is simply connected, we can introduce a potential $\psi$ such that the velocity ${\mathbf{u}}$ can be written ${\mathbf{u}}=(-\partial_{y}\psi,\partial_{x}\psi)^{\top}$ in such a way that the the first eigenvalue of the Stokes operator coincides with the first eigenvalue of the buckling problem [27, Chapter 11]. Our existence result is thus linked to the theorem of Ashbaugh & Bucur [1], which asserts the existence of an optimal domain for the buckling problem in the class of simply connected domains. We will comment more on this aspect of the problem in section 1.3.
@@
The local optimality of a shape will be investigated by means of Hadamard perturbations [30, Chapter 5]. This already requires some regularity of the shape under consideration. To be more precise: if ${\Omega}$ is a bounded domain with $\mathscr{C}^{2}$ boundary, if $\Phi\in W^{3,\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ we define, for any $t\in(-1;1)$ small enough,
@@
This definition is meaningful as, for any $t$ small enough, $(\mathrm{Id}+t\Phi)$ is a smooth diffeomorphism. For a given shape functional $F$ and a given shape ${\Omega}$, we say that $F$ is differentiable at ${\Omega}$ if, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$, the limit
@@
exists and if it is a linear form in $\Phi$. In this case, this limit is called the first-order shape derivative of $F$ at ${\Omega}$ in the direction $\Phi$.
@@
The first-order optimality conditions for (5) read: if ${\Omega}^{*}$ is a smooth optimal set for any smooth vector field $\Phi$, there holds
@@
We say that $F$ is twice-differentiable at ${\Omega}$ if it is differentiable and if, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ the limit
$\langle d^{2}F(\Omega){\Phi},{\Phi}\rangle=\lim_{t\to 0}\frac{F({\Omega}_{t\Phi})-F({\Omega})-\langle dF(\Omega),{\Phi}\rangle}{t^{2}}$
@@
    There exists a constant $c_{0}>0$ such that, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$,
@@
This theorem is to be expected as it is a generalisation of the classical Faber-Krahn inequality to the case of the Stokes operator.
@@
The differentiability of $\mathcal{F}$ at ${\mathbb{B}}_{2}$ is immediate if $\lambda_{1}({\mathbb{B}}_{2})$ is a simple eigenvalue. We refer to [30, Chapter 5], or to the implicit function theorem of Mignot, Murat & Puel [39] (for its application in the context of shape derivatives, see e.g. [19]).
@@
A much more surprising result is that the Faber-Krahn inequality is not true for the Stokes operator in $\mathbb{R}^{3}$. To the best of our knowledge, this is the first result of this kind, where the dimension has an influence on the optimality of the unit ball for the lowest eigenvalue of a differential operator. More precisely, we have:
@@
The first eigenvalue $\lambda_{1}({\mathbb{B}}_{3})$ has multiplicity 3. $\mathcal{F}$ is semi-differentiable at ${\mathbb{B}}_{2}$, but does not satisfy the first-order optimality conditions (6): there exists a vector field $\Phi$ such that
$\langle\partial\mathcal{F}({\mathbb{B}}_{3})\Phi,\Phi\rangle<0.$
@@
Let ${\Omega}^{*}$ be a solution of (5). If ${\Omega}^{*}$ has a $\mathscr{C}^{2}$ boundary then $\lambda_{1}({\Omega}^{*})$ is a simple eigenvalue. Furthermore, if ${\mathbf{u}}$ is an associated first eigenfunction, then $\|({\nabla}{\mathbf{u}})\nu\|$ is constant on $\partial{\Omega}$.
@@
The eigenvalues of the Stokes operator are generically simple with respect to the domain; we refer to Ortega & Zuazua and Chitour, Kateb & Long [17, 41]. Regarding the multiplicity of eigenvalues of the Stokes operator let us also mention the recent work of Falocchi & Gazzola [24]; it deals with Navier boundary conditions rather than with Dirichlet ones. Nevertheless, for an optimal domain (for another eigenvalue than the first one), multiplicity is very often expected [28, chapter 11].
@@
It is actually easy to prove that, at an optimiser, $\lambda_{1}({\Omega}^{*})$ has multiplicity at most 2. This will follow from considering the semi-differential of the eigenvalue. However, in order to lower the possible multiplicity to 1, we will draw inspiration from the work of Gerner [25, Proof of Theorem 2].
Although it is highly unlikely that these necessary optimality conditions will provide a full characterisation of the optimiser, the Poincaré-Hopf theorem (also known as the hairy ball theorem, see [16, Theorem 34.1] and [46, Theorem 39.7]) easily implies the following corollary:
@@
Assume ${\Omega}^{*}$ has a $\mathscr{C}^{3}$ boundary. If $\Sigma$ is a connected component of $\partial{\Omega}^{*}$, $\Sigma$ has Euler characteristic 1, and is thus homeomorphic to a torus.
Indeed, it suffices to observe that, as for any eigenfunction ${\mathbf{u}}$ in ${\Omega}$, as ${\mathbf{u}}=0\,,{\nabla}\cdot{\mathbf{u}}=0$ in ${\Omega}$, the vector field $({\nabla}{\mathbf{u}})\nu$ is tangential to $\partial{\Omega}$ (see Remark 5). Thus, Theorem 4 yields the applicability of the hairy ball theorem. It is notable that this line of reasoning was previously used in [21, 22].
@@
Indeed, writing, in coordinates, $\nu=\left(\nu_{i}\right)_{1\leqslant i\leqslant d}$ and ${\mathbf{u}}_{\Omega}=\left(u_{i}\right)_{1\leqslant i\leqslant d}$, we have
$\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,\nu\rangle=\sum_{i=1}^{d}\nu_{i}\left(\sum_{j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\nu_{j}\right)\quad\text{a.e. on }\partial{\Omega}.$
Since ${\mathbf{u}}_{\Omega}\equiv 0$ on $\partial{\Omega}$ there holds
$\forall i,j,\quad\frac{\partial u_{i}}{\partial x_{j}}\nu_{i}=\frac{\partial u_{j}}{\partial x_{i}}\nu_{j}\quad\text{a.e. on }\partial{\Omega}.$
@@
$\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,\nu\rangle$ $=\sum_{i=1}^{d}\nu_{i}\left(\sum_{j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\nu_{j}\right)=\sum_{i=1}^{d}\frac{\partial u_{i}}{\partial x_{i}}\left(\sum_{j=1}^{d}\nu_{j}^{2}\right)={\nabla}\cdot{\mathbf{u}}=0$

### v2

@@ Abstract
Keywords:
Shape derivation, Shape semi-differentiation, Spectral optimisation, Stokes operator.
AMS Classification:
47A10, 49Q10, 76D07.
@@ 1 Introduction
In this article, we focus on a spectral optimisation problem governed by the Stokes operator; the latter is crucial in the analysis of fluid motions. In order to motivate this question, let us observe that the eigenvalues of the Stokes operator can be physically interpreted as the decay frequencies of the eigenmodes of a fluid. Each associated eigenmode represents a specific fluid motion. The eigenvalues are also related to the characteristic time scales of the dynamics of the fluid. Furthermore, it should be noted that the eigenvalues of the Stokes operator also appear naturally when dealing with the long-time behaviour of solutions to the (nonlinear) Navier-Stokes equations [2, 26]. Our main goal here will thus be to further the understanding of the influence of the geometry of the domain on the first Stokes eigenvalue.
@@ 1.1 Setting
The optimisation of spectral quantities with respect to a domain is a central question in shape optimisation and in the calculus of variations. Throughout the article, $\Omega$ will denote a subset of the ambient space $\mathbb{R}^{d}$, with $d\in\{2,3\}$. The specific choice of dimension $d$ will be specified for each result. Of particular importance is the paradigmatic question of minimising (with respect to the domain ${\Omega}$) the first eigenvalue of a differential operator under a volume constraint on ${\Omega}$. The celebrated Faber-Krahn inequality asserts that, when the operator is the (scalar) Dirichlet-Laplacian, the ball minimises the first eigenvalue with a volume constraint. In the present paper, we investigate the minimisation of the first eigenvalue of the Dirichlet-Stokes operator. Namely, let ${\Omega}$ be a bounded, smooth open set, and let $W^{1,2}_{0}({\Omega};\mathbb{R}^{d})$ be defined as the usual Sobolev space (here, the functions are $\mathbb{R}^{d}$-valued; see also the Notations paragraph below). Consider the first eigenvalue
$\lambda_{1}({\Omega}):=\min_{\begin{subarray}{c}{\mathbf{u}}\in W^{1,2}_{0}({\Omega};\mathbb{R}^{d})\\ \nabla\cdot{\mathbf{u}}=0\text{ in }{\Omega},\ {\mathbf{u}}\neq 0\end{subarray}}\frac{\int_{\Omega}\|{\nabla}{\mathbf{u}}\|^{2}}{\int_{\Omega}\|{\mathbf{u}}\|^{2}},$ (1)
@@
${\nabla}{\mathbf{u}}=\left(\frac{\partial u_{i}}{\partial x_{j}}\right)_{1\leqslant i,j\leqslant d}.$
[…]
$\|\nabla{\mathbf{u}}\|^{2}={\nabla}{\mathbf{u}}\colon{\nabla}{\mathbf{u}}=\sum_{i,j=1}^{d}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{2}.$
This eigenvalue is associated with the eigen-equation
$\begin{cases}-\Delta{\mathbf{u}}+{\nabla}p=\lambda_{1}({\Omega}){\mathbf{u}}&\text{ in }{\Omega}\,,\\ {\nabla}\cdot{\mathbf{u}}=0&\text{ in }{\Omega}\,,\\ {\mathbf{u}}=0&\text{ on }\partial{\Omega}.\end{cases}$ (2)
In (2) the function $p$ (which is unique up to an additive constant) is the pressure associated with ${\mathbf{u}}$ , which can be interpreted as the Lagrange multiplier attached to the incompressibility constraint ${\nabla}\cdot{\mathbf{u}}=0$.
@@
The constant $V_{0}>0$ is a given volume constraint (which will be immaterial as we will work with a scale invariant formulation of (3)).
@@
    First, we obtain an existence result in the class of quasi-open sets (this is the natural framework for such existence results; we refer to Definition 1) This is Theorem 1. The method of proof relies on the concentration-compactness principle of Lions [44], which was adapted by Bucur [3] to the setting of shape optimisation. The main difficulty here is to handle the incompressibility condition.
@@
We discuss related works in details in section 1.3 but let us already highlight some aspects of our problem. One of the first results in spectral shape optimisation is the celebrated Faber-Krahn inequality [23, 38], which asserts that, under a volume constraint, the ball minimises the first eigenvalue of the Dirichlet-Laplacian in any dimension. This inequality can be derived through numerous methods, some of which can be generalised to the minimisation of other spectral quantities; the ball is very often the minimiser (or at least a local minimiser) for the first eigenvalue of several scalar differential operators. Thus, Theorem 3 hints at a deeper discrepancy between the scalar and the vectorial case. In general, let us observe that the literature devoted to spectral optimisation problems in the vectorial case is scarce. To the best of our knowledge, the results closest to ours were derived very recently by Enciso, Gerner & Peralta-Salas [21, 22, 28] in the case of the ${\mathbf{curl}}$ operator, a problem which was also investigated by Cantarella, DeTurck, Gluck & Teytel in the early 2000’s [11].
@@
    For any $k\in\mathbb{N}\,,p\in[1;+\infty)$, $W^{k,p}({\Omega})$ (resp. $W^{k,p}_{0}({\Omega})$) denotes the usual Sobolev space of order $k$ and index $p$ (resp. the functions of $W^{k,p}({\Omega})$ whose trace on $\partial{\Omega}$ is zero). Likewise, $W^{k,p}({\Omega};\mathbb{R}^{d})$ (resp. $W^{k,p}_{0}({\Omega};\mathbb{R}^{d})$) denotes the Sobolev space of order $k$ and index $p$ of vector-valued functions, each coordinate of which is a $W^{k,p}({\Omega})$ (resp. $W^{k,p}_{0}({\Omega})$) function.
@@
    The double-dot product of two matrices $A=(a_{ij})_{1\leqslant i,j\leqslant N}$ and $B=(b_{ij})_{1\leqslant i,j\leqslant N}$ is the real number $A\colon B$ given by $A\colon B=\sum_{i,j=1}^{N}a_{ij}b_{ij}$.
@@
    $\mathbb{B}_{d}=\{x\in\mathbb{R}^{d}\ |\ \|x\|<1\}$ is the Euclidean unit ball of $\mathbb{R}^{d}$.
@@
    $\|\mathbf{X}\|$ denotes either the Euclidean norm of $\mathbb{R}^{d}$ if $\mathbf{X}$ is a vector, or the Frobenius norm of $\mathbf{X}$ if $\mathbf{X}$ is a square matrix of size $d$; in other words $\|\mathbf{X}\|^{2}=\mathbf{X}:\mathbf{X}$.
  8. 8.
@@
  9. 9.
    $\operatorname{Per}(\Omega)$ denotes the De Giorgi perimeter of $\Omega$.
@@
To set our problem in an appropriate (from the point of view of existence properties) framework, we recall the definition of quasi-open sets [34, Chapter 3]:
@@
$\mathrm{cap}(\omega):=\sup_{\begin{subarray}{c}K\textnormal{compact}\\ K\subset\omega\end{subarray}}\,\inf_{\begin{subarray}{c}v\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{d})\\ v\geqslant 1\text{ on }K\end{subarray}}\int_{\mathbb{R}^{d}}\left|{\nabla}v\right|^{2}.$
@@
Quasi-open sets are the natural framework in which to consider spectral shape optimisation problems [10, 57]. For any given ${\Omega}\in\mathscr{O}$ we can define [34, Chapter 3] the Sobolev space $W^{1,2}_{0}({\Omega})$ as the set of functions $v\in W^{1,2}(\mathbb{R}^{d})$ that are equal to 0 on $\mathbb{R}^{d}\backslash{\Omega}$ up to a set of zero capacity. For any ${\Omega}\in\mathscr{O}$ we set
$\lambda_{1}({\Omega}):=\min_{\begin{subarray}{c}{\mathbf{u}}\in W^{1,2}_{0}({\Omega};\mathbb{R}^{3})\color[rgb]{0,0,0}\\ \nabla\cdot{\mathbf{u}}=0\text{ in }{\Omega}\,,{\mathbf{u}}\neq 0\end{subarray}}\frac{\int_{\Omega}\|{\nabla}{\mathbf{u}}\|^{2}}{\int_{\Omega}\|{\mathbf{u}}\|^{2}}.$ (4)
@@ 1.2 Main results of the paper
The proof of this theorem hinges, in this vectorial setting, on the approach to the Faber-Krahn inequality of Bucur & Freitas [5, Proposition 3.1] or Bucur & Varchon [6], which in turn relies on the concentration-compactness principle of Lions [44]. However, because of the incompressibility constraint, we have to modify the concentration-compactness approach, which in turns prohibits us from guaranteeing that ${\Omega}^{*}$ is bounded.
Of course, if $\mathscr{O}$ were to be replaced with $\mathscr{O}_{D}:=\{{\Omega}\in\mathscr{O}\,,\mathscr{O}\subset D\}$ for a given compact set $D$ (“box constraint”), the existence of a minimiser would follow from an adaptation of the Buttazzo-Dal Maso theorem [10], although one should be cautious when handling the zero-divergence constraint. For the sake of completeness, we include this result in Appendix B. We refer to section 1.3 for more references on shape optimisation without box constraints.
@@
When working in the two-dimensional case, if the domain ${\Omega}$ is simply connected, we can introduce a potential $\psi$ such that the velocity ${\mathbf{u}}$ can be written ${\mathbf{u}}=(-\partial_{y}\psi,\partial_{x}\psi)^{\top}$, so that the first eigenvalue of the Stokes operator coincides with the first eigenvalue of the buckling problem [31, Chapter 11]. The function $\psi$ is called the stream function; we refer for instance to [56, Lemma 2.5] for its existence and uniqueness. Its existence is a special case of the Helmoltz-Hodge decomposition in dimension 2. Our existence result is thus linked to the theorem of Ashbaugh & Bucur [1], which asserts the existence of an optimal domain for the buckling problem in the class of simply connected domains. We will comment more on this aspect of the problem in section 1.3.
@@
The local optimality of a shape will be investigated by means of Hadamard perturbations [34, Chapter 5]. This already requires some regularity of the shape under consideration. To be more precise: if ${\Omega}$ is a bounded domain with $\mathscr{C}^{2}$ boundary, if $\Phi\in W^{3,\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$ we define, for any $t\in(-1;1)$ small enough,
@@
This definition is meaningful as, for any $t$ small enough, $(\mathrm{Id}+t\Phi)$ is a smooth diffeomorphism. For a given shape functional $F$ and a given shape ${\Omega}$, we say that $F$ is differentiable at ${\Omega}$ if, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$, the limit
@@
exists, and if it is a linear form in $\Phi$. In this case, this limit is called the first-order shape derivative of $F$ at ${\Omega}$ in the direction $\Phi$.
@@
whenever the limit above exists, and defines a linear form in $\Phi\in W^{3,\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$.
@@
The first-order optimality conditions for (5) read: if ${\Omega}^{*}$ is a smooth optimal set then, for any smooth vector field $\Phi$, there holds
@@
The first condition above, called the Euler inequality, can often be reduced to an equality depending on the nature of the admissible set of shapes. This will be made precise in what follows.
We say that $F$ is twice differentiable at ${\Omega}$ if it is differentiable and if, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$ the limit
$\langle d^{2}F(\Omega){\Phi},{\Phi}\rangle=\lim_{t\to 0}\frac{F({\Omega}_{t\Phi})-F({\Omega})-t\langle dF(\Omega),{\Phi}\rangle}{t^{2}}$
@@
    There exists a constant $c_{0}>0$ such that, for any $\Phi\in W^{3,\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $\langle\Phi,\nu\rangle\in\{1,\cos(\cdot),\sin(\cdot)\}^{\perp}$, where $\perp$ denotes the orthogonal for the $L^{2}(\partial{\mathbb{B}})$ inner product,
@@
This theorem is to be expected as it is a generalisation of the classical Faber-Krahn inequality to the case of the Stokes operator. The condition $\langle\Phi,\nu\rangle\in\{1,\cos,\sin\}^{\perp}$ reflects the dilation and translation invariance of the functional $\mathcal{F}$.
@@
The differentiability of $\mathcal{F}$ at ${\mathbb{B}}_{2}$ is immediate if $\lambda_{1}({\mathbb{B}}_{2})$ is a simple eigenvalue. We refer to [34, Chapter 5], or to the implicit function theorem of Mignot, Murat & Puel [46] (for its application in the context of shape derivatives, see e.g. [19]).
Remark 3.
If a minimiser $\Omega^{*}$ of (5) in dimension 2 is smooth and simply connected, then $\Omega^{*}$ is necessarily a disc. The proof is due to N.B. Willms and H.F. Weinberger. They did not publish it, but one can find it in the paper of B. Kawohl in [35]. We also refer to [31, Section 11.3.4]. This result is proved by establishing an equivalence between the minimisation of the first eigenvalue of the Stokes-Dirichlet operator and the minimisation of the solution of the buckling problem. Detailed explanations of the buckling problem and the equivalence can be found in section 1.3. Once this equivalence is established, the proof that a regular and simply connected minimiser is a disc results from the study of an overdetermined problem.
Finally, let us point to the recent article [27], which was brought to our attention during the revision of this paper: an interesting discussion is given, linking several overdetermined problems including the one mentioned above.
@@
A much more surprising result is that the Faber-Krahn inequality is not true for the Stokes operator in $\mathbb{R}^{3}$. To the best of our knowledge, this is the first published result of this kind, where the dimension has an influence on the optimality of the unit ball for the lowest eigenvalue of a differential operator.
Remark 4.
Shortly after uploading a preprint of this paper, W. Gerner has informed us that similar (unpublished) results [29] had been obtained for some related spectral optimisation problems in the context of Riemaniann geometry, where the ball is optimal in the two-dimensional case, but not in the three-dimensional one. In the euclidean setting, the examples he gave were of a Stokes-like operator, with the additional boundary condition $\langle{\mathbf{curl}}({\mathbf{u}}),\nu\rangle=0$ on $\partial{\Omega}$, as well as the requirement that ${\mathbf{u}}$ be orthogonal to harmonic fields (that is, ${\mathrm{div}}$ and ${\mathbf{curl}}$ free fields).
More precisely, we have:
@@
The first eigenvalue $\lambda_{1}({\mathbb{B}}_{3})$ has multiplicity 3. $\mathcal{F}$ is semi-differentiable at ${\mathbb{B}}_{3}$, but does not satisfy the first-order optimality conditions (6): there exists a vector field $\Phi\in W^{3,\infty}(\mathbb{R}^{3};\mathbb{R}^{3})$ such that
$\langle\partial\mathcal{F}({\mathbb{B}}_{3}),\Phi\rangle<0.$
@@
Let ${\Omega}^{*}$ be a solution of (5) with $d=3$. If ${\Omega}^{*}$ has a $\mathscr{C}^{2,\alpha}$ boundary with $\alpha\in(0,1)$ then $\lambda_{1}({\Omega}^{*})$ is a simple eigenvalue. Furthermore, if ${\mathbf{u}}$ is an associated first eigenfunction, then $\|({\nabla}{\mathbf{u}})\nu\|$ is constant on $\partial{\Omega}$.
@@
The eigenvalues of the Stokes operator are generically simple with respect to the domain; we refer to Ortega & Zuazua and Chitour, Kateb & Long [17, 48]. Regarding the multiplicity of eigenvalues of the Stokes operator let us also mention the recent work of Falocchi & Gazzola [24, 25]; it deals with Navier boundary conditions rather than with Dirichlet ones. Nevertheless, for an optimal domain (for an eigenvalue other than the first one), multiplicity is very often expected [32, chapter 11].
@@
It is actually easy to prove that, at an optimiser, $\lambda_{1}({\Omega}^{*})$ has multiplicity at most 2. This will follow from considering the semi-differential of the eigenvalue. However, in order to lower the possible multiplicity to 1, we draw inspiration from the work of Gerner [28, Proof of Theorem 2].
Although it is highly unlikely that these necessary optimality conditions will provide a full characterisation of the optimiser, the Poincaré-Hopf theorem (also known as the hairy ball theorem, see [16, Theorem 34.1] and [53, Theorem 39.7]) easily implies the following corollary:
@@
Assume ${\Omega}^{*}$ has a $\mathscr{C}^{2,\alpha}$ boundary with $\alpha\in(0,1)$. If $\Sigma$ is a connected component of $\partial{\Omega}^{*}$, $\Sigma$ has Euler characteristic 0, and is thus homeomorphic to a torus.
Indeed, it suffices to observe that, as for any eigenfunction ${\mathbf{u}}$ in ${\Omega}$, as ${\mathbf{u}}=0$ in $\partial{\Omega}$ and ${\nabla}\cdot{\mathbf{u}}=0$ in ${\Omega}$, the vector field $({\nabla}{\mathbf{u}})\nu$ is tangential to $\partial{\Omega}$ (see Remark 7). Thus, Theorem 4 yields the applicability of the hairy ball theorem. It is notable that this line of reasoning was previously used in [21, 22]. These results are in good agreement with the numerical simulations of [43], which sets out to study numerical approximations of spectral optimisation problems for the Stokes operator.
@@
Indeed, as ${\mathbf{u}}_{\Omega}=0$ on $\partial{\Omega}$, we have, for any $1\leqslant i\leqslant d$,
${\nabla}u_{i}=\frac{\partial u_{i}}{\partial\nu}\nu,$
whence, for any $1\leqslant i,j\leqslant d$,
$\frac{\partial u_{i}}{\partial x_{j}}=\frac{\partial u_{i}}{\partial\nu}{\nu_{j}}.$
@@
$\langle({\nabla}{\mathbf{u}}_{\Omega})\nu,\nu\rangle$ $=$ $\sum_{i=1}^{d}\nu_{i}\left(\sum_{j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\nu_{j}\right)=\sum_{i=1}^{d}\frac{\partial u_{i}}{\partial\nu}\nu_{i}\left(\sum_{j=1}^{d}\nu_{j}^{2}\right)$
$=$ $\sum_{i=1}^{d}\frac{\partial u_{i}}{\partial\nu}\nu_{i}=\sum_{i=1}^{d}\frac{\partial u_{i}}{\partial x_{i}}={\nabla}\cdot{\mathbf{u}}=0$